### v2

# SC-LoRA: Balancing Efficient Fine-tuning and Knowledge Preservation via Subspace-Constrained LoRA

Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), are indispensable for efficiently customizing Large Language Models (LLMs). However, vanilla LoRA suffers from slow convergence speed and knowledge forgetting problems. Recent studies have leveraged the power of designed LoRA initialization, to enhance the fine-tuning efficiency, or to preserve knowledge in the pre-trained LLM. However, none of these works can address the two cases at the same time. To this end, we introduce Subspace-Constrained LoRA (SC-LoRA), a novel LoRA initialization framework engineered to navigate the trade-off between efficient fine-tuning and knowledge preservation. We achieve this by constraining the output of trainable LoRA adapters in a low-rank subspace, where the context information of fine-tuning data is most preserved while the context information of preserved knowledge is least retained, in a balanced way. Such constraint enables the trainable weights to primarily focus on the main features of fine-tuning data while avoiding damaging the preserved knowledge features. We provide theoretical analysis on our method, and conduct extensive experiments including safety preservation and world knowledge preservation, on various downstream tasks. In our experiments, SC-LoRA succeeds in delivering superior fine-tuning performance while markedly diminishing knowledge forgetting, surpassing contemporary LoRA initialization methods.
SC-LoRA: Balancing Efficient Fine-tuning and Knowledge Preservation via Subspace-Constrained LoRA
Minrui Luo∗1,2, Fuhang Kuang∗2, Yu Wang3, Zirui Liu2, Tianxing He†2,1thanks: Equal contribution.thanks: Corresponding author. 1Shanghai Qi Zhi Institute 2Institute for Interdisciplinary Information Sciences, Tsinghua University 3Institute of Information Engineering, Chinese Academy of Sciences {luomr22,kfh22,liu-zr22}@mails.tsinghua.edu.cn; wangyu2002@iie.ac.cnhetianxing@mail.tsinghua.edu.cn
1 Introduction
Fine-tuning effectively adapts large language models to downstream tasks (Luo et al., 2025; Yu et al., 2024). Due to the high computational cost of full fine-tuning, parameter-efficient fine-tuning (PEFT) methods (Xu et al., 2023; Han et al., 2024) have been proposed to reduce the number of trainable parameters while maintaining good fine-tuning performance. Among various PEFT methods, LoRA (Hu et al., 2022) is a simple yet efficient approach that introduces trainable low-rank adaptation modules for tuning. While LoRA offers significant parameter efficiency, it has two important problems: (1) the convergence speed of the fine-tuning process is relatively slow due to the noise and zero initialization of adapter modules; (2) it potentially leads to catastrophic forgetting problem (Goodfellow et al., 2015) as other fine-tuning methods do, such as harming the world knowledge stored in pre-trained LLMs (Yang et al., 2024), and degrading the safety of aligned LLMs (Qi et al., 2024).
Recent works have found that carefully designed initialization on LoRA adapters can solve these problems. (Meng et al., 2024) initializes LoRA adapters by parts of Singular Value Decomposition (SVD) of original weight $W_{0}$, leading to faster convergence and improved performance by encapsulating the most significant information stored in $W_{0}$. Later works (Yang et al., 2024; Paischer et al., 2024) initialize LoRA weights based on semantic information stored in the activations of each layer on the target fine-tuning dataset. These data-driven approaches successfully enhance the fine-tuning speed and performance. Towards catastrophic forgetting problem in LoRA fine-tuning, (Yang et al., 2024) proposes to initialize LoRA weights by the least principal directions of world knowledge data features, successfully alleviating the forgetting problem. However, these works can only solve either side of the two problems, but do not consider the trade-off between enhancing fine-tuning performance and preserving pre-trained knowledge, which is a common need when doing parameter-efficient fine-tuning.
In this paper, we introduce Subspace-Constrained LoRA, a balanced LoRA scheme that achieves both better fine-tuning results and good preservation of knowledge in LLMs. Specifically, we compute directions of linear layer output that align with the principal directions of fine-tuning data and at the same time are orthogonal to the principal directions of preserved knowledge. These directions are then used to initialize the adapter weights, constraining the output vectors (of each adapter layer) in a subspace spanned by these directions. This constraint intuitively makes the updating terms to focus on the fine-tuning data information, while avoiding affecting the preserved knowledge. By extensive experiments, we verify that by such constraint on balanced directions, our method achieves both efficient fine-tuning and excellent knowledge preservation, solving the problems that previous methods cannot address. In conclusion, our contribution includes:
  1. 1.
    We propose SC-LoRA, a balanced LoRA scheme that can achieve efficient fine-tuning and knowledge preservation at the same time, which previous methods cannot handle.
  2. 2.
    We provide theoretical proofs to explain our strategies, including analysis on subspace selection and initialization setting.
  3. 3.
    We conduct extensive experiments regarding both safety preservation and world knowledge preservation on various downstream tasks, verifying the effectiveness of our method.
2 Related Work
Parameter-Efficient Fine-Tuning (PEFT).
Modern large language models (LLMs) with billions of parameters face significant computational and memory challenges during full-parameter fine-tuning on downstream tasks, motivating the development of Parameter-Efficient Fine-Tuning (PEFT) methods that optimize only a small amount of parameters while maintaining model performance (Xu et al., 2023; Han et al., 2024).
Common PEFT approaches include partial fine-tuning (Ben Zaken et al., 2022; Bu et al., 2024) that only tune part of the parameters; soft prompt finetuning (Hambardzumyan et al., 2021; Lester et al., 2021), where trainable prompts are appended to inputs with model parameters frozen; adapter tuning (Houlsby et al., 2019; Lin et al., 2020; Rücklé et al., 2021; Karimi Mahabadi et al., 2021; Pfeiffer et al., 2021; He et al., 2022; Wang et al., 2022; Lei et al., 2023) which inserts additional trainable layers into LLMs and fix the base model parameters; and LoRA(Hu et al., 2022; Aghajanyan et al., 2021), which decomposes weight updates into low-rank matrices. Different from other approaches, LoRA does not change the original model architecture or incurring extra computational cost during inference since the extra adapters can be merged into original parameters.
LoRA Initialization.
Multiple LoRA initialization methods have been proposed, with the aim of improving training efficiency or obtaining other abilities.
PiSSA (Meng et al., 2024) argued that the default initialization of “Gaussian noise (He et al., 2015) and zero” to the adapters can lead to slow convergence. Hence they propose to apply singular value decomposition to original weight matrices and utilizes the top components to initialize LoRA, encapsulating the most significant information stored in original weights. CorDA (Yang et al., 2024) utilizes covariance matrices of data context, and takes the first (or last) singular vectors after context-oriented decomposition as initialization of LoRA adapters. They propose two different modes, one for improving fine-tuning performance and the other for mitigating world knowledge forgetting. Similar to CorDA, EVA (Paischer et al., 2024) feeds fine-tuning data into the model, applies sigular value decomposition to activation covariance matrices, and takes top singular vectors as initialization weights. LoRA-GA (Wang et al., 2024b) also utilizes data context but applies decomposition on the gradient. (Hayou et al., 2024) analyze the initialization of LoRA adapters, and have shown how the asymmetry of two low rank matrices affects training dynamics.
Harmful Finetuning Attack and Defense strategies.
To prevent potential misuse, LLMs usually undergo specific training to align them with human values before deployment (Ouyang et al., 2022; Bai et al., 2022). Nevertheless, jailbreak attacks employ carefully designed inputs to circumvent this alignment, with prominent methods including Greedy Coordinate Gradient (GCG) (Zou et al., 2023), AutoDAN (Liu et al., 2023), and PAIR (Chao et al., 2023). Beyond these direct attacks, fine-tuning can also undermine a model’s safety alignment, even when non-harmful data is used (Qi et al., 2024; He et al., 2024).
Consequently, researchers have developed various defense strategies against such fine-tuning risks, generally falling into following approaches: enhancing the original safety alignment (Huang et al., 2024c, a), restricting the scope of fine-tuning parameters (Wei et al., 2024; Li et al., 2025), mixing additional safety data (Wang et al., 2024a; Huang et al., 2024b), modifying the loss function (Qi et al., 2025) and post-fine-tuning processing (Yia et al., 2024; Hsu et al., 2024). Different from previous works, our method focuses on an alternative approach of mitigating safety risks during fine-tuning by modifying initialization and model architecture, without mixing safety data during fine-tuning or appending prefix during inference time - both of which would incur computation overhead.
World Knowledge Forgetting.
Catastrophic forgetting (McCloskey and Cohen, 1989) is a phenomenon when models lose previously acquired knowledge when adapting to new tasks, and has been extensively studied in deep learning. Early approaches to solve the problem include knowledge distillation (Li and Hoiem, 2018), rehearsal (Riemer et al., 2019) and dynamic architectures (Yan et al., 2021). For large language models, preserving world knowledge remains challenging due to massive pre-training data and model size. Recent efforts mitigate forgetting by freezing pre-trained layers while introducing new adapters (Wu et al., 2024; Dou et al., 2024). Recently (Yang et al., 2024) proposed CorDA with Knowledge-Preserved Adaptation (KPA) mode, addressing world knowledge forgetting through LoRA initialization.
3 Method
Below, we first review the vanilla LoRA, and describe our proposed SC-LoRA method.
((a)) LoRA
((b)) SC-LoRA
Figure 1: Comparison of LoRA with default Kaiming initialization and our proposed SC-LoRA. (a) LoRA initializes down-projection matrix A by Gaussian noise and up-projection matrix B by zero matrix. (b) Our SC-LoRA initializes A by $Q_{r}^{\top}W_{0}$ and B by $Q_{r}$, where $Q_{r}$ consists of $r$ orthonormal vectors as columns obtained by Algorithm 1.
3.1 LoRA
Following the hypothesis that the update of weight matrices presents a low rank structure (Aghajanyan et al., 2021), LoRA (Hu et al., 2022) uses the product of two trainable low-rank matrices to learn the weight change while keeping the original weight matrices frozen. To express in mathematical form, LoRA adds low-rank adapters $A,B$ to original weight matrix $W_{0}$ by $W^{\prime}=W_{0}+BA$, where $W^{\prime},W_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, $A\in\mathbb{R}^{r\times d_{\mathrm{in}}}$, $B\in\mathbb{R}^{d_{\mathrm{out}}\times r}$, $r\ll\min(d_{\mathrm{in}},d_{\mathrm{out}})$. When fine-tuning, $W_{0}$ is kept frozen, and $A,B$ are trainable parameters.
From the default initialization scheme of LoRA, $A$ is initialized by Kaiming Initialization (He et al., 2015) while $B$ is initialized by zero matrix. Consequently, the adapter term $BA=O$ and $W^{\prime}=W_{0}$ at the start of fine-tuning, ensuring the coherence with the model before fine-tuning. For initializations with non-zero adapter $BA$ (Meng et al., 2024; Yang et al., 2024; Wang et al., 2024b), the frozen weights are adjusted to the residual term $W_{\mathrm{res}}=W_{0}-B_{\mathrm{init}}A_{\mathrm{init}}$. Then the adapted weight is $W^{\prime}=W_{0}-B_{\mathrm{init}}A_{\mathrm{init}}+BA=W_{\mathrm{res}}+BA$. In transformer-based LLMs, LoRA adapters are applied to weight matrices within the self-attention and multilayer perceptron (MLP) layers.
3.2 SC-LoRA
Known as catastrophic forgetting problem Chen et al. (2020), a large language model often performs worse on its pre-trained knowledge after fine-tuning on a downstream task. To this end, we consider fine-tuning a large language model on downstream task $T_{+}$, while preserving its ability on the other task $T_{-}$. Consider the output of a linear layer $h=W_{0}x=W_{\mathrm{res}}x+B_{\mathrm{init}}A_{\mathrm{init}}x$. We denote $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ the distribution of $h$ when the model is fed with data from $T_{+}$ and $T_{-}$, respectively. Our aim is to initialize $A,B$ within the $r$-rank constraint so that $BAx$ preserves the most of $\mathcal{P}_{+}$ and the least of $\mathcal{P}_{-}$, so that after initialization, the trainable term $BAx$ is constrained to primarily focus on $\mathcal{P}_{+}$ while avoiding modifying $\mathcal{P}_{-}$. This is equivalent to identify a low-dimensional subspace $S\subset\mathbb{R}^{d_{\mathrm{out}}}$ with rank $r$, on which the projection of $\mathcal{P}_{+}$ is mostly preserved and the projection of $\mathcal{P}_{-}$ is mostly eliminated. To evaluate such property of subspace $S$, we define the following reward:
Definition 1.
For a subspace $S\subset\mathbb{R}^{d_{\mathrm{out}}}$ of dimension $r$, define the reward $R(S)$ over $\mathcal{P}_{\pm}$ as:
$R(S)$ $=(1-\beta)\mathbb{E}_{X_{+}\sim\mathcal{P}_{+}}\left[\left\|\Pi_{S}\left(X_{+}\right)\right\|_{2}^{2}\right]$ (1)
$-\beta\mathbb{E}_{X_{-}\sim\mathcal{P}_{-}}\left[\left\|\Pi_{S}(X_{-})\right\|_{2}^{2}\right],$
where $\beta\in[0,1]$ is a hyperparameter to tune. Here $\Pi_{S}:\mathbb{R}^{d_{\mathrm{out}}}\to S$ denote the orthogonal projection operator onto $S$. See Appendix A.1 for mathematical definition of $\Pi_{S}$.
The first term of $R(S)$ quantifies the context information of $T_{+}$ contained in subspace $S$, while the second penalizes that of $T_{-}$. We use $\beta$ to balance the trade-off between focusing on $T_{+}$ and preservation on $T_{-}$. Given the objective to maximize $R(S)$, in the following we provide Theorem 1 to compute the optimal subspace and then use it to set our LoRA initialization scheme.
Theorem 1.
Let $\mathrm{Cov}_{+},\mathrm{Cov}_{-}$ be the covariance matrices of random vectors $X_{+}\sim\mathcal{P}_{+}$ and $X_{-}\sim\mathcal{P}_{-}$, respectively:
$\mathrm{Cov}_{+}$ $=\mathbb{E}\left[X_{+}X_{+}^{\top}\right],$ (2)
$\mathrm{Cov}_{-}$ $=\mathbb{E}\left[X_{-}X_{-}^{\top}\right].$ (3)
And let
$\Delta\mathrm{Cov}=(1-\beta)\mathrm{Cov}_{+}-\beta\mathrm{Cov}_{-}.$ (4)
Then do eigenvalue decomposition of $\Delta\mathrm{Cov}$ and take the first $r$ eigenvectors $\{q_{i}\}_{i\in[r]}$ with the largest eigenvalues. Then, if following condition holds, the reward $R(S)$ is maximized:
$S=\mathrm{span}\left(\left\{q_{i}\right\}_{i\in[r]}\right).$ (5)
Proof.
See Appendix A.2. ∎
Theorem 1 shows the steps to compute the optimal subspace that maximized $R(S)$. Then, to constrain the updating output term $BAx$ in the subspace $S$, we propose our LoRA initialization method:
$B_{\mathrm{init}}$ $=(q_{1}\,q_{2}\,\cdots\,q_{r}),$ (6)
$A_{\mathrm{init}}$ $=(q_{1}\,q_{2}\,\cdots\,q_{r})^{\top}W_{0},$ (7)
$W_{\mathrm{res}}$ $=W_{0}-B_{\mathrm{init}}A_{\mathrm{init}},$ (8)
as illustrated in Figure 1(b). To explain the initialization setting, we provide the following theorem:
Theorem 2.
Let $h,x$ be the output and input of the original linear layer $W_{0}$, satisfying $h=W_{0}x$. When $A,B$ are initialized by Equations 7, 8, the following property holds:
$B_{\mathrm{init}}A_{\mathrm{init}}x=\Pi_{S}(h)\in S,\,\forall x\in\mathbb{R}^{d_{\mathrm{in}}}.$ (9)
Proof.
See Appendix A.3. ∎
Together with Theorem 1, our initialization method has the following properties: When $\beta=0$ and the model is fed with data from task $T_{+}$, $h$ follows distribution $\mathcal{P}_{+}$, then the norm of the updating term $BAx$ is maximized, providing the most context information of $T_{+}$ for training; When $\beta=1$ and the model is fed with data from task $T_{-}$, $h$ follows distribution $\mathcal{P}_{-}$, then the norm of $BAx$ is minimized, passing the least context information of $T_{-}$ to trainable parameters. When $\beta\in(0,1)$, it is the balance between the two cases. The property indicates that, during fine-tuning, the trainable weights are updating more on features related to $T_{+}$ and less on features related to $T_{-}$, and hence enhancing learning $T_{+}$ while avoiding damaging information related to $T_{-}$.
The pseudo-code of our initialization algorithm is shown in Algorithm 1. In practice, it is hard to format the true distribution and covariance of output vectors, so we approximate them by feeding hundreds of samples into the model, and use the collection of output vectors to approximate the distribution.
Algorithm 1 SC-LoRA initialization.
0:  Datasets $\mathcal{D}_{+},\mathcal{D}_{-}$ from tasks $T_{+},T_{-}$, respectively.
1:  Let $B_{+}=|\mathcal{D}_{+}|,B_{-}=|\mathcal{D}_{-}|$.
2:  Separately feed samples in $\mathcal{D}_{+},\mathcal{D}_{-}$ into the pre-trained model, collect batched output $\hat{X}_{+}\in\mathbb{R}^{d_{out}\times B_{+}},\hat{X}_{-}\in\mathbb{R}^{d_{out}\times B_{-}}$ of each linear layer. Within each sample, the output vector is summed over all tokens.
3:  $\mathrm{Cov}_{+}\leftarrow\frac{1}{B_{+}}\hat{X}_{+}\hat{X}_{+}^{\top}$.
4:  $\mathrm{Cov}_{-}\leftarrow\frac{1}{B_{-}}\hat{X}_{-}\hat{X}_{-}^{\top}$.
5:  Do eigenvalue decomposition on $\Delta\mathrm{Cov}=(1-\beta)\mathrm{Cov}_{+}-\beta\mathrm{Cov}_{-}$, and take the first $r$ eigenvectors $\{q_{i}\}_{i\in r}$ with the largest eigenvalues.
6:  $Q_{r}\leftarrow(q_{1}\,q_{2}\,\cdots\,q_{r})$.
7:  $B_{\mathrm{init}}\leftarrow Q_{r}$.
8:  $A_{\mathrm{init}}\leftarrow Q_{r}^{\top}W_{0}$
9:  $W_{\mathrm{res}}\leftarrow W_{0}-B_{\mathrm{init}}A_{\mathrm{init}}$.
4 Experiments
In the experiments below, we compare SC-LoRA with 5 baselines:
(1) Full fine-tuning. Fine-tune on all parameters of the model;
(2) Vanilla LoRA (Hu et al., 2022). Fine-tune only on LoRA adapters, with $B$ initialized with Gaussian noise (He et al., 2015), and $A$ initialized by zero;
(3) PiSSA (Meng et al., 2024), for efficient fine-tuning. It applies SVD on pre-trained weight $W_{0}$ and initializes LoRA adapters by the main parts of decomposition;
(4) CorDA (Yang et al., 2024) Instruction-Previewed Adaptation (IPA) mode, for efficient fine-tuning. It feeds fine-tuning data into the model to get the covariance of activations, applies self-defined context-oriented decomposition, and initializes LoRA adapters with principal directions obtained in decomposition;
(5) CorDA Knowledge-Preserved Adaptation (KPA) mode, for knowledge preservation. The initialization algorithm is basically the same as IPA mode except that it feeds preserved knowledge data and take the least principal directions for initialization.
For initialization of CorDA IPA and KPA mode, we calculate the covariance matrices with 256 samples from fine-tuning dataset and preserved knowledge dataset, respectively with 256 samples. We use AdamW optimizer Loshchilov and Hutter (2019) with the following hyper-parameters: batch size 128, learning rate 2e-5 (except for experiment in Section 4.2, where we tune the learning rate of baselines for better performance), cosine annealing learning rate schedule, warm-up ratio 0.03, and no weight decay. The rank of LoRA and its variants are all set to 128 for comparison. For SC-LoRA, we tune the hyperparameter $\beta$ to find a good balanced result. All experiment results are obtained by running on only one seed.
Below we discuss results in three settings: (1) Preservation of safety when fine-tuning on benign data; (2) Preservation of safety when fine-tuning on poisoned data; (3) Preservation of world knowledge when fine-tuning on math task.
[…]
Llama-2-7b-Chat - 1.100 1.212 24.13
Full fine-tuning 6738M 1.364 5.455 51.41
LoRA 320M 1.176 2.424 50.32
PiSSA 320M 1.252 4.242 51.87
CorDA IPA 320M 1.209 3.333 44.61
CorDA KPA 320M 1.106 0.606 50.89
SC-LoRA $\beta=0.5$ 320M 1.161 1.818 52.54
$\beta=0.7$ 320M 1.148 1.818 52.07
$\beta=0.9$ 320M 1.097 0.000 51.67
Table 1: Results of Safety preservation and fine-tuning performance when training on benign dataset Samsum. #Params is the number of trainable parameters. HS and HR denote harmfulness score and harmfulness rate respectively.
4.1 Safety Preservation on Benign Finetuning
(Qi et al., 2024) has shown that fine-tuning on benign data can compromise the safety of aligned LLMs. In this setting, we aim to preserve the safety of aligned LLM while providing efficient fine-tuning on downstream tasks. Following the experimental settings by (Qi et al., 2025), we fine-tune Llama-2-7b-Chat model with safety alignment (Touvron et al., 2023) on Samsum (Gliwa et al., 2019) for 1 epoch. Samsum is a dataset for conversation summarization task, containing 14732 training samples.
To initialize our SC-LoRA model, we randomly select 256 samples from fine-tuning dataset ($\mathcal{D}_{+}$) Samsum to compute covariance matrix $\mathrm{Cov}_{+}$ for each linear layer, then use 256 harmful-question&refusal-answer pairs (as the safety dataset $\mathcal{D}_{-}$) provided by (Qi et al., 2025) to compute $\mathrm{Cov}_{-}$. These two collections of samples are also used to compute the covariance matrices of CorDA IPA and CorDA KPA respectively.
For utility evaluation, we employ the standard ROUGE-1 score (Lin, 2004) for Samsum. For safety evaluation, we let the fine-tuned models to generate answers for 330 malicious questions provided by (Qi et al., 2024) (distinct from malicious questions for initialization) and employ DeepSeek-V3 (DeepSeek-AI et al., 2025) API to judge the harmfulness, assigning each answer an integer score from 1 (safe) to 5 (most harmful). We report the average score as harmfulness score of the model and the fraction of maximum-risk responses (score = 5) as harmfulness rate. Lower values for both metrics indicate stronger safety of the model.
As shown in Table 1, SC-LoRA achieves high utility, even surpassing full fine-tuning on Samsum dataset when $\beta=0.5$. At the same time, SC-LoRA shows almost no safety degradation compared to the model before fine-tuning, while all baselines except CorDA KPA present notable safety degradation, since they are not designed for knowledge preservation. However, the utilities of all fine-tuning methods (except for CorDA IPA) are generally close. We hypothesize that the task of summarization is quite simple, so training for only 1 epoch is enough for utility convergence. Also, the results of SC-LoRA shows that when $\beta$ is increasing, the safety preservation becomes better while utility is decreasing. This aligns with our design of $\beta$ to balance the trade-off.
[…]
Llama-2-7b-Chat - 1.100 1.212 -
Full fine-tuning 6738M 2.248 23.94 41.47
LoRA lr=2e-5 320M 1.118 1.212 31.69
lr=5e-5 320M 2.276 23.64 37.68
lr=1e-4 320M 3.155 41.52 41.93
PiSSA 320M 2.379 29.39 41.77
CorDA IPA 320M 4.239 67.27 43.75
CorDA KPA 320M 1.127 1.212 40.33
SC-LoRA $\beta=0.5$ 320M 1.630 10.91 45.56
$\beta=0.7$ 320M 1.224 3.030 45.26
$\beta=0.9$ 320M 1.136 1.212 45.26
Table 2: Results of safety preservation and fine-tuning performance when training on poisoned dataset MetaMathQA with 1% malicious question-answer pairs.
4.2 Safety Preservation on Data Poisoning Attack
Harmful data injection is a common attack method to degrade the safety of LLMs during fine-tuning (Huang et al., 2024a, b, c). In this experiment, we aim to preserve safety in poisoned data scenarios. To construct the poisoned dataset, we first take 25600 data samples of MetaMathQA (Yu et al., 2024), then replace 1% of the data by harmful question-answer pairs provided by (Qi et al., 2024). We train each method for 1 epoch on the poisoned dataset. We use 256 samples from MetaMathQA for the initialization of SC-LoRA and CorDA IPA. The safety samples used for the initialization of SC-LoRA and CorDA KPA are the same with the previous experiment (Section 4.1).
For utility evaluation, we compute the answer accuracy on the validation set of GSM8k (Cobbe et al., 2021). Safety evaluation follows the setting in the previous section 4.1. For better comparability, we tune the learning rate of LoRA to 2e-5, 5e-5 and 1e-4. The learning rate for other methods is fixed to 2e-5.
From the results in Table 2, we can observe that the data points exhibit a wider spread among these methods, both in utility and safety metric. Compared to the original model, SC-LoRA ($\beta=0.9$) exhibits almost no safety degradation, and achieves best utility, even surpassing full fine-tuning by 3.79 points. When increasing the learning rate, LoRA shows a sharp decline in safety alignment while math ability is increasing. LoRA (lr=2e-5) and CorDA KPA, though preserving safety well, are insufficient in fine-tuning performance compared to our method. PiSSA and CorDA IPA, though showing their capacity in better fine-tuning, heavily degrades the safety of the model. This again shows the potential of our method to enhance the utility of the model and preserve safety at the same time, even when the fine-tuning dataset contains a small fraction of harmful content. Also, the utility and safety of SC-LoRA follows the same trend as in fine-tuning on benign data when $\beta$ is increasing, supporting the sedign of our method.
[…]
Llama-2-7b - 52.52 18.86 5.86 25.75 - - -
Full fine-tuning 6738M 47.42 4.16 6.64 19.41 50.27 6.94 28.60
LoRA 320M 46.81 1.05 7.04 18.30 41.77 5.46 23.62
PiSSA 320M 47.44 3.32 6.84 19.20 51.63 7.70 29.67
CorDA IPA 320M 30.20 9.83 5.41 15.15 51.40 8.34 29.87
CorDA KPA 320M 46.21 10.64 7.33 21.39 45.03 6.54 25.79
SC-LoRA $\beta=0$ 320M 44.26 5.18 7.19 18.88 53.53 8.98 31.25
$\beta=0.5$ 320M 48.91 7.70 6.89 21.17 53.37 8.62 31.00
$\beta=0.8$ 320M 50.52 10.64 7.04 22.73 52.46 7.62 30.04
Table 3: Results of world knowledge preservation and math ability after fine-tuning on MetaMATH.
4.3 World Knowledge Preservation
Pre-trained LLMs also have other pre-trained knowledge that is easy to lose after fine-tuning on downstream tasks, such as world knowledge Yang et al. (2024). In this setting, we aim to preserve the intrinsic world knowledge (e.g., common sense) within the pre-trained LLM while providing efficient fine-tuning on downstream tasks. We fine-tune the Llama-2-7b model (Touvron et al., 2023) on math task and evaluate its math ability (utility) and world knowledge performance. We train on 100000 samples of MetaMathQA (Yu et al., 2024) for 1 epoch and evaluate its math ability on GSM8k (Cobbe et al., 2021) and MATH (Yu et al., 2024) validation sets. World knowledge is evaluated by the exact matching score on TriviaQA (Joshi et al., 2017), NQ-open (Lee et al., 2019), and WebQS (Berant et al., 2013) through Evaluation-Harness Gao et al. (2024). We select 256 random samples from NQ-open as world knowledge samples used for the initialization of SC-LoRA and CorDA KPA, and 256 random samples from MetaMathQA as fine-tuning dataset for initializing SC-LoRA and CorDA IPA.
As shown in Table 3, the results of full fine-tuning and LoRA show the degradation on world knowledge when fine-tuning on downstream task MetaMathQA. SC-LoRA achieves best math ability (surpassing full fine-tuning), and preserves world knowledge relatively well. When $\beta=0.8$, it surpasses all baselines on both utility and knowledge preservation. Also, from the results of SC-LoRA, we can see a clear trend when increasing $\beta$, that the knowledge preservation ability is increasing while the utility is decreasing, which aligns with our design methodology for $\beta$ in Section 3. More details will be shown in Section 4.4 to analyze this trend.
Figure 2: Relations between $\beta$ and knowledge preservation performance. The experiment setting of the left figure is described in Section 4.1, while that of the right figure is described in Section 4.3. Lower harmfulness score or higher world knowledge score indicates better performance on knowledge preservation.
Figure 3: Relations between $\beta$ and fine-tuning performance. The experiment setting of the left figure is described in Section 4.1, while that of the right figure is conducted in Section 4.3. The right figure shows clear monotonicity with $\beta$, while such trend does not occur in the left figure.
4.4 Experimental Analysis on the Functionality of Hyper-Parameter $\beta$
As explained in Section 3, the value of $\beta$ balance the trade-off between knowledge preservation and fine-tuning efficiency. Intuitively, when increasing $\beta$, there exists a trend that the fine-tuning performance will drop and the knowledge preservation ability will increase. While we have observed this trend in the previous section, we illustrate the trend more explicitly in Figure 2 and 3. In Figure 2, both two curves shows knowledge preservation improvement when $\beta$ is increasing: one for safety increasing, and the other for world knowledge preservation improvement. In Figure 3, the math ability decreases when $\beta$ is increasing, aligning with our expectations. The utility on Samsum, however, does not show evident trend as $\beta$ varies, but fluctuating around 0.52. We hypothesize that the task of summerization is quite simple, so whatever the value of $\beta$, it is sufficient for utility convergence during fine-tuning.
These trends give experimental support to our method design, that by adjusting $\beta$ we can balance the trade-off. Interestingly, a linear relationship was observed between $\beta$ values and knowledge preservation in some experimental settings.
5 Conclusion
Aimed to balance the trade-off between efficient fine-tuning and knowledge preservation, this paper presents a data-driven LoRA initialization that utilizes the subspace constrain, in order to strengthen the target knowledge while downgrading its influence on preserved knowledge. Theoretical analysis are provided to support our method, including the choice of subspace and the initialization setting. We conduct extensive experiments regrading safety preservation and world knowledge preservation, during fine-tuning on various downstream tasks such as math and summarization. The results of experiments strongly demonstrate that our method can not only promote fine-tuning performance on downstream tasks, but also preserve the intrinsic knowledge stored in pre-trained model, surpassing contemporary LoRA initialization methods.
6 Limitations
Although SC-LoRA can successfully handle both efficient fine-tuning and knowledge preservation at the same time, it still has drawbacks.
First, SC-LoRA is just a LoRA initialization method, and does not strongly constrain the updates during fine-tuning process. Hence after fine-tuning on more complex tasks and with more steps, the knowledge preservation ability can also drop (see the preservation drop of NQ-open in Table 3 for example).
Second, its application on preserving other types of knowledge remains unexplored. Future work may consider applying SC-LoRA to preserving multimodal large language model’s performance on pre-training tasks (Zhai et al., 2024) or large language model’s reasoning ability.
These aspects provide promising directions for future researches.
Acknowledgments
We are grateful to Prof. Jingzhao Zhang from Tsinghua University for helpful discussions and constructive comments on this work.
Appendix A Definitions and Proofs
A.1 Mathematical Definition of $\Pi_{S}$
Definition 2.
Suppose $S$ is a subspace of $\mathbb{R}^{n}$ of dimension $r$, and let $\{q_{i}\}_{i\in[r]}$ be an orthonormal basis of $S$, then the orthogonal projection operator onto $S$, denoted $\Pi_{S}$, is defined as:
$\Pi_{S}(x)$ $=\sum_{i=1}^{r}(q_{i}^{\top}x)q_{i}$ (10)
$=\sum_{i=1}^{r}(q_{i}q_{i}^{\top})x,\forall x\in\mathbb{R}^{n}.$
Note: the selection of the orthonormal basis does not affect $\Pi_{S}$.
A.2 Proof for Theorem 1
Proof.
Suppose for some subspace $S\subset\mathbb{R}^{d}$ (ignore the subscript of $d_{\mathrm{out}}$ for simplicity) of dimension $r$, there exists an orthonormal basis $\left\{v_{i}\right\}_{i\in[r]}$ that spans $S$, that is $S=\mathrm{span}\left(\left\{v_{i}\right\}_{i\in[r]}\right)$.
For simplicity, denote
$\tilde{I}_{r}=\sum_{i\in[r]}v_{i}v_{i}^{\top},$ (11)
then the following equality holds:
$\tilde{I}_{r}^{\top}\tilde{I}_{r}$ $=\sum_{i\in[r]}\sum_{j\in[r]}v_{i}v_{i}^{\top}v_{j}v_{j}^{\top}$ (12)
$=\sum_{i\in[r]}\sum_{j\in[r]}v_{i}\langle v_{i},v_{j}\rangle v_{j}^{\top}$
$=\sum_{i\in[r]}\sum_{j\in[r]}\delta_{ij}v_{i}v_{j}^{\top}$
$=\sum_{i\in[r]}v_{i}v_{i}^{\top}=\tilde{I}_{r}.$
From property of projection,
$\Pi_{S}\left(X_{\pm}\right)$ $=\sum_{i=1}^{r}\langle X_{\pm},v_{i}\rangle v_{i}=\sum_{i=1}^{r}v_{i}v_{i}^{\top}X_{\pm}$ (13)
$=\left(\sum_{i=1}^{r}v_{i}v_{i}^{\top}\right)X_{\pm}=\tilde{I}_{r}X_{\pm}.$
Thus
$\mathbb{E}_{X_{\pm}\sim\mathcal{P_{\pm}}}\left[\left\|\Pi_{S}\left(X_{\pm}\right)\right\|_{2}^{2}\right]$ (14)
$=$ $\mathbb{E}_{X_{\pm}\sim\mathcal{P_{\pm}}}\left[\left\|\tilde{I}_{r}X_{\pm}\right\|_{2}^{2}\right]$
$=$ $\mathbb{E}_{X_{\pm}\sim\mathcal{P}_{\pm}}\left[\mathrm{tr}\left(X_{\pm}^{\top}\tilde{I}_{r}^{\top}\tilde{I}_{r}X_{\pm}\right)\right]$
$=$ $\mathbb{E}_{X_{\pm}\sim\mathcal{P}_{\pm}}\left[\mathrm{tr}\left(X_{\pm}^{\top}\tilde{I}_{r}X_{\pm}\right)\right]$
$=$ $\mathbb{E}_{X_{\pm}\sim\mathcal{P}_{\pm}}\left[\mathrm{tr}\left(\tilde{I}_{r}X_{\pm}X_{\pm}^{\top}\right)\right]$
$=$ $\mathrm{tr}\left(\tilde{I}_{r}\mathbb{E}_{X_{\pm}\sim\mathcal{P}_{\pm}}\left[X_{\pm}X_{\pm}^{\top}\right]\right)$
$=$ $\mathrm{tr}\left(\tilde{I}_{r}\mathrm{Cov}_{\pm}\right).$
Suppose the spectral decomposition of $(1-\beta)\mathrm{Cov}(X_{+})-\beta\mathrm{Cov}(X_{-})$ is $Q\Sigma Q^{\top}$, where $Q=(q_{1}\,q_{2}\,\cdots\,q_{d})$, $\Sigma$ is diagonal with eigenvalues sorted in descending order. Then we have
$R(S)$ $=(1-\beta)\mathbb{E}_{X_{+}\sim\mathcal{P}_{+}}\left[\left\|\Pi_{S}\left(X_{+}\right)\right\|_{2}^{2}\right]$ (15)
$-\beta\mathbb{E}_{X_{-}\sim\mathcal{P}_{-}}\left[\left\|\Pi_{S}(X_{-})\right\|_{2}^{2}\right]$
$=(1-\beta)\mathrm{tr}\left(\tilde{I}_{r}\mathrm{Cov}_{+}\right)-\beta\mathrm{tr}\left(\tilde{I}_{r}\mathrm{Cov}_{-}\right)$
$=\mathrm{tr}\left(\tilde{I}_{r}\Delta\mathrm{Cov}\right)$
$=\sum_{i\in[r]}\mathrm{tr}\left(v_{i}v_{i}^{\top}Q\Sigma Q^{\top}\right)$
$=\sum_{i\in[r]}v_{i}^{\top}Q\Sigma Q^{\top}v_{i}.$
Extend $\left\{v_{i}\right\}_{i\in[r]}$ to a complete orthonormal basis $\left\{v_{i}\right\}_{i\in[d]}$ for $\mathbb{R}^{d}$, and denote $u_{i}=Q_{i}^{\top}v_{i}$. Since $Q$ is an orthogonal matrix, $\{u_{i}\}_{i\in[d]}$ is also an orthonormal basis for $\mathbb{R}^{d}$. From Ky Fan’s theorem on eigenvalues, $\max\left(\sum_{i\in[r]}v_{i}^{\top}Q\Sigma Q^{\top}v_{i}\right)=\sum_{i\in[r]}\Sigma_{ii}$, and one can easily verify that the condition above achieves the maximum.
For the if and only if part (adding the condition of eigenvalue gap): suppose $U=(u_{1}\,u_{2}\,\cdots\,u_{d})^{\top}$ as an orthogonal matrix, then
$R(\{v_{i}\}_{i\in[r]})$ $=\sum_{i\in[r]}u_{i}^{\top}\Sigma u_{i}$ (16)
$=\sum_{i\in[r]}\sum_{j=1}^{d}\Sigma_{jj}U_{ij}^{2}$
$=\sum_{j=1}^{d}\left(\Sigma_{jj}\sum_{i\in[r]}U_{ij}^{2}\right).$
From property of orthogonal matrix, $\sum_{i\in[r]}U_{ij}^{2}\leq 1$ and $\sum_{j=1}^{d}\sum_{i\in[r]}U_{ij}^{2}=r$, then to maximize $R$, from the additional assumption we need $\sum_{i\in[r]}U_{ij}^{2}=\begin{cases}1,&1\leq j\leq r\\ 0,&r+1\leq j\leq d\end{cases}$, which is equivalent to
$U_{1:r,1:d}^{\top}U_{1:r,1:d}=\begin{pmatrix}I_{r}&O\\ O&O\end{pmatrix}.$ (17)
From $U_{1:r,1:d}=(v_{1}\,v_{2}\,\cdots\,v_{r})^{\top}Q$, we know that this is also equivalent to
$(v_{1}\,v_{2}\,\cdots\,v_{r})(v_{1}\,v_{2}\,\cdots\,v_{r})^{\top}=Q\begin{pmatrix}I_{r}&O\\ O&O\end{pmatrix}Q^{\top},$ (18)
which is also written as
$\sum_{i=1}^{r}v_{i}v_{i}^{\top}=\sum_{i=1}^{r}q_{i}q_{i}^{\top}.$ (19)
Indicating $S=\mathrm{span}\left(\left\{q_{i}\right\}_{i\in[r]}\right)$.
A.3 Proof of Theorem 2
Proof.
Denote $Q_{r}=(q_{1}\,q_{2}\,\cdots\,q_{r})$.
Since $\{q_{i}\}_{i\in[r]}$ is a orthonormal basis that spans $S$, from definition of orthogonal projection we have
$\Pi_{S}(h)=\sum_{i=1}^{r}q_{i}q_{i}^{\top}h=Q_{r}Q_{r}^{\top}h.$ (20)
Thus $\forall x\in\mathbb{R}^{d_{\mathrm{in}}}$, we have
$B_{\mathrm{init}}A_{\mathrm{init}}x=Q_{r}Q_{r}^{\top}W_{0}x=Q_{r}Q_{r}^{\top}h=\Pi_{S}(h),$ (21)
which completes the proof.
Appendix B Numerical instability in sparse sample setting
When the sample size is much larger than the output activation dimension, $\min(|\mathcal{D}_{+}|,|\mathcal{D}_{-}|)\gg d_{\mathrm{out}}$, setting $\beta\in[0,1]$ causes no issue. However, when samples are sparse (specifically, when the number of negative-task sample $N_{-}<d_{\mathrm{out}}-r$, setting $\beta=1$ introduces multiple valid solutions in the spectral decomposition step due to high-dimensional freedom in the null space of $\mathrm{Cov}_{-}$. Mathematically, the rank of $\mathrm{Cov}_{-}$ is at most $N_{-}$, resulting in a null space of dimension $d_{\mathrm{out}}-\mathrm{rank}(\mathrm{Cov}_{-})\geq d_{\mathrm{out}}-N_{-}>r$. Consequently, any arbitrary set of $r$ orthonormal vectors in this null space can satisfy the decomposition criterion, leading to non-unique initialization of parameters $A$ and $B$.
To mitigate this instability, we recommend setting $1-\beta$ to a small positive value (rather than exactly zero). This retains the regularization from $\mathrm{Cov}_{+}$ in the objective function, which constrains the null space ambiguity and stabilizes the spectral decomposition, empirically improves fine-tuning performance.